how to slove this 2x+4<10
Answer by sdmmadam@yahoo.com(530)

2x + 4 < 10
2x < 10 - 4 (adding -4 to both the sides of the inequality or subracting 4 from both the sides of the inequality)
2x < 6
which is on dividing by 2>0
x < 3
Note: Division or multiplication by a positive quantity does not alter the inequality.
Verification:
Give x = 3 which implies 2x + 4 = 10 which is not less than 10
Give x = something slightly more than 3 say x = 3.1 which gives
2x + 4 = 2x3.1 +4 = 6.2 + 4 = 10.2 which is once again not less than 10
Thus x<3 alone will work.